in a traingle one angle is twice the measure of the second angle the third angle is three times the measure of the second angle .find the measure of each angle
Answer by checkley77(12844)

2x, x, 3x are the 3 angles.
2x=x+3x=180
6x=180
x=180/6
x=30 ans. for the second angle
2*30=60 ans. for the first angle.
3*30=90
ans. for the third angle.
Proof:
60+30+90=180
180=180